OpenStudy (anonymous):

how do solve logarithmic equations? 3 log 2x=4

OpenStudy (anonymous):

Remember \[\log_{10} x = y\] is 10^y = x log 2x =4/3 10^(4/3) = 2x x= (10^(4/3))/2